OpenStudy (anonymous):

Explain the difference between using the cosine ratio to solve for a missing angle in a right triangle versus using the secant ratio. You must use complete sentences and any evidence needed (such as an example) to prove your point of view.

OpenStudy (tkhunny):

You make the complete sentences. Extremely little difference. \(\cos(\alpha) = \dfrac{1}{\sec(\alpha)}\) It uses the SAMe information.

OpenStudy (anonymous):

Thank you :) @tkhunny